Run Coordination Meetings:Fall2017 Spring2018 Run/March recovery

From GlueXWiki
Jump to: navigation, search

Runplan for restarting the Spring 2018 run after the March down time due to the transformer failure

Time chart for the recovery

Before the beam is established, check that:

  • Tagger quad is on at -4.2A
  • PS magnet is on at 910A
  • Sweeping magnet is on at 190A
  • Collimator is in blocking position
  • All radiators are retracted

Once the beam position is good on 5C11B (see numbers on the left white board)

  • Fill the LH2 target
  • Insert the 2×10-5 RL Al. radiator (remember to call MCC and ask to mask the FSD each time we insert/remove something from the electron beam line)
  • Insert the profiler
  • Request 50 nA. Check that the beam position is at x=0,y=0 on the profiler

Check the radiation monitor levels (see table on the bottom right of the white board) at 50 nA with and without the 2×10-5 RL Al. radiator

Do Harp scans (ours and MCC).

If the result of one of the tasks above does not look good, ask MCC to improve the beam tune.

Tell/email Hovanes that he can check the beam convergence (we do not need this result to proceed).

Insert 1×10-4 RL Al. radiator and request 100 nA Put the 5mm collimator in. Check that the beam position on the Act. Col. corresponds to the expected numbers (see left white board) when beam is centered on the profiler.

Retract profiler

PS rates check:

  • Retract the 1×10-4 RL Al. radiator
  • Insert the production diamond (JD70-100)
  • Insert the TPol 75 μm Be converter
  • Request 150 nA

High granularity PS scaler coincidence rate should be around 2.7±0.3 kHz. (see HDLOG entry #3541345 for expected rates at 150 nA). If PS rate on diamond is not right, then:

  • Insert the 4.5×10-4RL Al. radiator (the one in the goniometer)
  • Request 150 nA

High granularity PS scaler coincidence rate should be about 1.4±0.2 kHz. If PS rate on Al. radiator is wrong too, call Hovanes to do the photon beam transmission optimization (i.e. collimator scan). This should take about 2h. If the PS rate on Al. radiator is good while the one on the diamond was not, email Hovanes that there might be a problem with the diamond alignment.

Once the PS rate is good, do:

  • CDC background studies (L. Penchev, 8h)
  • Standard production data until next day's morning, 7:30am
  • Harp scans (7:30am-8am)
  • Converter thickness study (A. Somov, 2h)
  • L1 trigger and SSP studies (A. Somov, 2h)
  • Diamond (JD70-100) scan to find a new spot (Hovanes, 2h). This is planned for the afternoon, so if Sacha completes his studies early, we can take some production data until a whatever time is convenient for the diamond work.
  • Take production data and daily harp scans until Monday morning.